3.38 g of AgNO3 react with BaCl2 to produce 1.43 g of AgCl according to the following balanced equation: 2AgNO3 + BaCl2 → 2 AgCl + Ba(NO3)2 What is the percent yield of AgCl in this reaction?
1
Expert's answer
2020-06-25T07:51:01-0400

According to the equation, moles AgNO3 = moles AgCl.

So the theoretical yield of this reaction can be calculated as:


mtheor(AgCl) = (m(AgNO3) / M(AgNO3)) x M (AgCl) = (3.38g / 169.87g/mol) x 143.32g/mol = 2.85 g.


%yield = (mactual / mtheor) x 100% = (1.43g / 2.85g) x 100% = 50.2%


Answer: 50.2%